SIMULTANEOUS ESTIMATION OF SEVERAL STRATUM MEANS UNDER ERROR-IN-VARIABLES SUPERPOPULATION MODELS

GUOHUA ZOU^{1} AND ALAN T. K. WAN^{2}

1Institute of Systems Science, Chinese Academy of Sciences, Beijing 100080, China 2Department of Management Sciences, City University of Hong Kong, Hong Kong, China

(Received January 5, 1998; revised November 4, 1998)

Abstract. This paper considers simultaneous estimation of means from several strata under error-in-variables superpopulation models. Necessary and sufficient conditions for an estimator to be admissible in the class of linear estimators are obtained.

Key words and phrases: Error-in-variables superpopulation models, mean vector of strata, admissibility, linear estimator.

——–